OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. inches Part B How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles

OpenStudy (anonymous):

I can heelp you

OpenStudy (igreen):

Set up a proportion. \(\dfrac{2}{25} = \dfrac{x}{60}\) Multiply 60 to both sides: \(\dfrac{2 \times 60}{25}\) Simplify that

OpenStudy (anonymous):

that is great because i keep getting the answer wrong

OpenStudy (anonymous):

25mile−−−>2inch divide by 25 throughout 1mile−−−>225inch multiply by 60 throughout 60mile−−>60×225inch

OpenStudy (anonymous):

so they are saying multiply 60 and 2 divide by 25

OpenStudy (anonymous):

so do i multiply 60 by 2

OpenStudy (anonymous):

Yes and divide by 25

OpenStudy (anonymous):

4.8

OpenStudy (igreen):

Correct, that's our answer.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

ya Great work can u fan and medal me thxs

OpenStudy (anonymous):

yeah thanks for the help but i put 4.8 and the answer is wrong

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@saynabahmed13

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u medal and fan me

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@saynabahmed13 i need your help on part b

OpenStudy (anonymous):

@saynabahmed13 i need you help